# Jones’ conjecture for Halin graphs and a bit more

### v2

@@ 1 Introduction
A classical theorem of Erdős and Pósa [4] says that if a graph $G$ does not contain more than $k$ pairwise disjoint cycles, then we can delete $f(k)\leq ck\log k$ vertices to make the graph acyclic. Jones’ conjecture [5] raises the same problem in planar graphs. It states that if a planar graph contains no more than $k$ pairwise disjoint cycles, then there are $2k$ vertices such that removing them from the graph gives a forest. Jones’ conjecture was proved for outerplanar graphs [5] and subcubic planar graphs [1]. Currently, the best known general result is that if the maximum number of disjoint cycles in an arbitrary planar graph is $k$, then there are $3k$ vertices such that removing them from the graph we get a forest, as proved independently by Chen et al. [3], Ma et al. [6] and Chappell et al. [2].
@@
In the following proof a good triangle is a triangle $xyz$, for which $d(x)=d(y)=3$ so that the vertices $x$ and $y$ are on the outer face. A good triangle for a vertex $u$ is a good triangle that does not contain $u$.
@@
If the degree of each vertex in a based planar graph $G$ is at least three and the degree of each vertex on the outer face is $3$, then for an arbitrary vertex $u$ on the outer face, there is a good triangle $xyz$ in the graph $G$ for the vertex $u$.
@@
If this forest is a tree, but it is not a star, then let us take a longest path in it. Consider one of the two penultimate vertices of the path (they are different because the graph is not a star). At least two neighbors of the penultimate vertex of this path will be leaves, since its degree is at least three, and apart from its neighbor on the path, it cannot have a neighbor that is not a leaf, because then it would not be a longest path. Two of these leaves will be adjacent on the outer face, so we have found a good triangle. We can do the same at the other end of the longest path. So, we found two vertex-disjoint good triangles. It follows that there is a good triangle for every vertex.
Consider the case where the resulting forest consists of several components. Let us choose one of the trees as we want. Its leaves divide the border of the outer face into arcs. Let us choose another tree; its vertices are on one of the resulting arcs. We may assume that neither tree of the two has only two vertices. If this were the case, then we can find suitable trees on both arcs formed by the two-point tree, and we can choose them. Denote $a$ and $b$ the two leaves of the first tree, which are the endpoints of the arc that contains the second tree, and denote $c$ and $d$ the two extreme leaves of the second tree on the arc so that the vertices follow each other in order $a,c,d,b$ at the border of the outer face. Denote $\mathcal{C}$ the arc $ab$ that contains the vertices $c$ and $d$. Delete the edges connected to $a$ and $c$ on the arc $ac$ contained in $\mathcal{C}$ (this edges are may be the same) and delete the edges connected to $b$ and $d$ on the arc $bd$ contained in $\mathcal{C}$ (this edges are may be the same) and connect with an edge the vertices $a$ and $b$ and the vertices $c$ and $d$. Thus, we get at least two components. We can use induction for two of them: $G_{1}$ that contains the vertices $a$ and $b$ and $G_{2}$ that contains the vertices $c$ and $d$. By induction, there is a good triangle for $a$ in $G_{1}$ and a good triangle for $c$ in $G_{2}$. None of the two triangles can use one of the two new edge, so they are vertex-disjoint good triangles in the original graph $G$. It follows that there is a good triangle for every vertex. (It is not necessary, but it is sufficient to find two disjoint triangles whose vertices on the outer face have degree $3$.) ∎
@@
If the degree of each vertex in a based planar graph $G$ is at least three, then for an arbitrary vertex $u$ on the outer face, there is a good triangle $xyz$ in the graph $G$ for the vertex $u$.
@@
Let us consider a vertex $v$ on the outer face whose degree is greater than three and what is not a cut vertex. Let us split the vertex $v$ into a vertex $v_{1}$ and a vertex $v_{2}$ ($v_{1}$ and $v_{2}$ should be adjacent). The two edges that are on the border of the outer face and run into the vertex $v$, after that let them run into the vertex $v_{1}$. The other edges that run into the vertex $v$, after that let them run into the vertex $v_{2}$. This operation maintains that the degree of each vertex is at least three (the degree of the vertices $v_{1}$ and $v_{2}$ is at least three and the degrees of the other vertices do not change). It is preserved that the graph is a planar graph and that every face is adjacent to the outer face. The value of $k$ has decreased for the graph, so we can apply the induction assumption to it. The vertex $v_{1}$ will not be included in a triangle. If the vertex $v_{2}$ is included in a good triangle for some vertex $u$, it corresponds to a good triangle in the original graph for $u$ in which the vertex $v$ is present instead of the vertex $v_{2}$.
The only problem is that the induction does not guarantee a good triangle for the vertex $v$. A Since for this we have to stipulate that the vertex $v_{2}$ is not included in the triangle, but the vertex $v_{2}$ is not on the border of the outer face. Assume by contradiction that every good triangle that has two three-degrees vertices on the outer face contains the vertex $v$. There must be at least two such triangles, since for every vertex different from the vertex $v$ we proved by induction that there is a good triangle that does not contain that vertex. Let these two good triangles be $vab$ and $vcd$ such that they are located in the order $vabcdv$ on the border of the outer face. $d(a)=d(b)=d(c)=d(d)=3$, so neither $a$ nor $d$ can be the neighbor of $v$ on the border of the outer face. The vertices $a$ and $d$ divide the boundary of the outer face into two arcs. Let us take the one that contains the vertices $b$ and $c$. This arc, together with the vertices $a,d$, and $v$, defines a cycle. Delete all the vertices that are inside the cycle (in the planar embedding) and all internal points of the arc and join the vertices $a$ and $d$ with an edge. The resulting graph satisfies the conditions of the lemma: the degree of every vertex is at least three (because $a$ and $d$ are adjacent to their neighbors on the border of the outer face, to $v$ and to each other) and every face is adjacent to the outer face. Furthermore, the number of vertices on the outer face whose degree is greater than three cannot increase. So we can use the induction assumption. In the resulting graph, we find a good triangle that does not contain $v$, this is also a suitable triangle in the original graph, so we have reached a contradiction. Since it cannot happen that we do not find a good triangle for the vertex $v$, the induction works and the statement of the lemma is true. ∎
@@
If there is a vertex $u$ in the graph $G$ of which degree is smaller than $2$, then we leave one of these vertex. Let the resulting graph be $G^{\prime}$. $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, since $u$ is not in any cycle. The statement is true for the graph $G^{\prime}$ by induction, so it is also true for the graph $G$.
@@
If there is a triangle $xyz$ in the graph $G$ such that $d(x)=2$, then let $G^{\prime}$ be the graph obtained by omitting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})\leq\textrm{cp}(G)-1$ since let we take a cycle packing in the graph $G^{\prime}$, in the graph $G$ we can also add the triangle $xyz$ to this cycle packing. Thus for every cycle packing in the graph $G^{\prime}$ we find a larger cycle packing in the graph $G$, so the maximum size of a cycle packing of $G$ necessarily bigger than the maximum size of a cycle packing of $G^{\prime}$. Let $\textrm{cp}(G)=k$ then $\textrm{cp}(G^{\prime})\leq k-1$, so $\textrm{fvs}(G^{\prime})\leq 2k-2$ by induction. Let be $F^{\prime}$ an arbitrary minimum feedback vertex set of the graph $G^{\prime}$, then $F=F^{\prime}\cup\{y,z\}$ is a feedback vertex set of the graph $G$ because the difference between the graph $G\setminus F$ and the graph $G^{\prime}\setminus F^{\prime}$ is only an isolated vertex. Thus, $\textrm{fvs}(G)\leq\textrm{fvs}(G^{\prime})+2\leq 2k$.
@@
If there is a vertex $u$ of degree two in the graph $G$, which is not in a triangle, then let $G^{\prime}$ be the graph obtained by omitting the vertex $u$ from the graph $G$ and connecting the two neighbors of the vertex $u$ with an edge. This graph is trivially still a based planar graph. All minimum feedback vertex set in this graph is also a minimum feedback vertex set in the original graph. If a minimum feedback vertex set in $G$ does not contain $u$, then it is a minimum feedback vertex set in $G^{\prime}$, too. If a minimum feedback vertex set in $G$ contains $u$, then let us delete $u$ from it and add a neighbor of $u$. So, we get a minimum feedback vertex set with the same size in $G^{\prime}$. Furthermore, similar is true for cycle packings: they are essentially identical in the two graphs. So $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, thus, by induction, we are done. (The statement is true for the graph $G^{\prime}$, so it is also true for the graph $G$.)
@@
Now, we may assume that the degree of each vertex in the graph is at least three. Then, according to Lemma 2.3, there is a triangle $xyz$, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face. Let $G^{\prime}$ be the graph obtained by deleting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})<\textrm{cp}(G)$, since the vertex $x$ cannot be included in any cycle in the graph $G^{\prime}$, so we can add the triangle $xyz$ to an arbitrary cycle packing of the graph $G^{\prime}$ and thus get a bigger cycle packing. Then let us take a minimum feedback vertex set of the graph $G^{\prime}$, adding the vertices $y$ and $z$ to this, we get a feedback vertex set of the graph $G$, whose size is at most $2\textrm{cp}(G^{\prime})+2\leq 2\cdot\textrm{cp}(G)$. ∎
We actually also proved the following stronger statement:
Corollary 2.5.
@@ 3 Concluding remarks
Problem 3.1.
Let $G$ be a planar graph in which there is a cycle which has at least one edge in common with every face. Prove, that $\textrm{fvs}(G)\leq 2\cdot\textrm{cp}(G)$.
@@
Let be $G$ a hamiltonian planar graph. Prove, that $\textrm{fvs}(G)\leq 2\cdot\textrm{cp}(G)$.

### v3

@@ 1 Introduction
A classical theorem of Erdős and Pósa [4] says that if a graph $G$ does not contain more than $k$ pairwise disjoint cycles, then we can delete $f(k)\leq ck\log k$ vertices to make the graph acyclic. Jones’ conjecture [5] raises the same problem in planar graphs. It states that if a planar graph contains no more than $k$ pairwise disjoint cycles, then there are $2k$ vertices such that removing them from the graph gives a forest. Jones’ conjecture was proved for outerplanar graphs [5] and subcubic planar graphs [1]. Currently, the best known general result is that if the maximum number of disjoint cycles in an arbitrary planar graph is $k$, then there are $3k$ vertices such that removing them from the graph we get a forest, as independently proved by Chen et al. [3], Ma et al. [6] and Chappell et al. [2].
@@
In the following proof a good triangle is a triangle xyzxyz, for which $d(x)=d(y)=3$ so that the vertices $x$ and $y$ are on the outer face. A good triangle for a vertex $u$ is a good triangle that does not contain $u$.
@@
If the degree of each vertex in a based planar graph $G$ is at least three and the degree of each vertex on the outer face is $3$, then for an arbitrary vertex $u$ on the outer face, there is a good triangle xyzxyz in the graph $G$ for the vertex $u$.
@@
If this forest is a tree, but it is not a star, then let us take a longest path in it. Consider one of the two penultimate vertices of the path (they are different because the graph is not a star). Every neighbor of that vertex, other than its neighbor on the path, is a leaf, because if not, then the path would not be a longest path. Thus the penultimate vertex has at least two leaf neighbors. Let us take two of them. The two leaf-edges divide the inner region into two subregions, the nonterminal vertices of the longest path are in one of these subregions. Now consider the other subregion. In this subregion every edge incident to the penultimate vertex is a leaf edge. (If not then either there is a longer path or the forest would not be connected.) So, every vertex of this subregion except the penultimate vertex is a leaf. Two of these leaves will be adjacent on the border of the outer face, so we have found a good triangle. We can do the same at the other end of the longest path. So, we found two vertex-disjoint good triangles. It follows that there is a good triangle for every vertex.
Consider the case where the resulting forest consists of several components. Let us choose one of the trees as we want. Its leaves divide the border of the outer face into arcs. Let us choose another tree; its vertices are on one of the resulting arcs. We prove that we may assume that neither of the two trees has only two vertices.
Assume that one of these trees has only two vertices. (If there are two, take any of them.) Then these two vertices divide the cycle into two arcs. No multiple edges, so there are additional vertices (and trees) on both arcs. Instead of the previously chosen trees, choose the tree in each arc that covers a smallest section of this arc. If $F$ is such a tree then no other tree is associated to the arc covered by $F$. So, $F$ should have more than 2 vertices as we have seen above.
Denote $a$ and $b$ the two leaves of the first tree, which are the endpoints of the associated arc that disjoint to the second tree, and denote $c$ and $d$ the two extreme leaves of the second tree on the arc so that the vertices follow each other in order $a,c,d,b$ at the border of the outer face. Denote $\mathcal{C}$ the arc $ab$ that contains the vertices $c$ and $d$. Delete the edges connected to $a$ and $c$ on the arc $ac$ contained in $\mathcal{C}$ (these edges are maybe the same) and delete the edges connected to $b$ and $d$ on the arc $bd$ contained in $\mathcal{C}$ (these edges are maybe the same) and connect with an edge the vertices $a$ and $b$ and the vertices $c$ and $d$. Thus, we get at least two components. We can use induction for two of them: $G_{1}$ that contains the vertices $a$ and $b$ and $G_{2}$ that contains the vertices $c$ and $d$. By induction, there is a good triangle for $a$ in $G_{1}$ and a good triangle for $c$ in $G_{2}$. Neither of the two triangles can use either of the two new edges, so they are vertex-disjoint good triangles in the original graph $G$. It follows that there is a good triangle for every vertex. (It is not necessary, but it is sufficient to find two disjoint triangles whose vertices on the outer face have degree $3$.) ∎
@@
If the degree of each vertex in a based planar graph $G$ is at least three, then for an arbitrary vertex $u$ on the outer face, there is a good triangle xyzxyz in the graph $G$ for the vertex $u$.
@@
Let us consider a vertex $v$ on the outer face whose degree is greater than three and that is not a cut vertex. Let us split the vertex $v$ into a vertex $v_{1}$ and a vertex $v_{2}$ ($v_{1}$ and $v_{2}$ should be adjacent). The two edges that are on the border of the outer face and run into the vertex $v$, after that let them run into the vertex $v_{1}$. The other edges that run into the vertex $v$, after that let them run into the vertex $v_{2}$. This operation maintains that the degree of each vertex is at least three (the degree of the vertices $v_{1}$ and $v_{2}$ is at least three and the degrees of the other vertices do not change). It is preserved that the graph is a planar graph and that every face is adjacent to the outer face. The value of $k$ has decreased for the graph, so we can apply the induction assumption to it. The vertex $v_{1}$ will not be included in a triangle. If the vertex $v_{2}$ is included in a good triangle for some vertex $u$, it corresponds to a good triangle in the original graph for $u$ in which the vertex $v$ is present instead of the vertex $v_{2}$.
The only problem is that the induction does not guarantee a good triangle for the vertex $v$. For this we have to stipulate that the vertex $v_{2}$ is not included in the triangle, but the vertex $v_{2}$ is not on the border of the outer face. Assume by contradiction that every good triangle that has two three-degrees vertices on the outer face contains the vertex $v$. There must be at least two such triangles, since for every vertex different from the vertex $v$ we proved by induction that there is a good triangle that does not contain that vertex. Let these two good triangles be vabvab and vcdvcd such that they are located in the order vabcdvvabcdv on the border of the outer face. $d(a)=d(b)=d(c)=d(d)=3$, so neither $a$ nor $d$ can be the neighbor of $v$ on the border of the outer face. The vertices $a$ and $d$ divide the boundary of the outer face into two arcs. Let us take the one that contains the vertices $b$ and $c$. This arc, together with the vertices $a,d$, and $v$, defines a cycle. Delete all the vertices that are inside the cycle (in the planar embedding) and all internal points of the arc and join the vertices $a$ and $d$ with an edge. The resulting graph satisfies the conditions of the lemma: the degree of every vertex is at least three (because $a$ and $d$ are adjacent to their neighbors on the border of the outer face, to $v$ and to each other) and every face is adjacent to the outer face. Furthermore, the number of vertices on the outer face whose degree is greater than three cannot increase. So we can use the induction assumption. In the resulting graph, we find a good triangle that does not contain $v$, this is also a suitable triangle in the original graph, so we have reached a contradiction. Since it cannot happen that we do not find a good triangle for the vertex $v$, the induction works and the statement of the lemma is true. ∎
@@
If there is a vertex $u$ in the graph $G$ of which degree is smaller than $2$, then we delete one of those vertices. Let the resulting graph be $G^{\prime}$. $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, since $u$ is not in any cycle. The statement is true for the graph $G^{\prime}$ by induction, so it is also true for the graph $G$.
@@
If there is a triangle xyzxyz in the graph $G$ such that $d(x)=2$, then let $G^{\prime}$ be the graph obtained by omitting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})\leq\textrm{cp}(G)-1$ since let we take a cycle packing in the graph $G^{\prime}$, in the graph $G$ we can also add the triangle xyzxyz to this cycle packing. Thus for every cycle packing in the graph $G^{\prime}$ we find a larger cycle packing in the graph $G$, so the maximum size of a cycle packing of $G$ is necessarily bigger than the maximum size of a cycle packing of $G^{\prime}$. Let $\textrm{cp}(G)=k$ then $\textrm{cp}(G^{\prime})\leq k-1$, so $\textrm{fvs}(G^{\prime})\leq 2k-2$ by induction. Let $F^{\prime}$ be an arbitrary minimum feedback vertex set of the graph $G^{\prime}$, then $F=F^{\prime}\cup\{y,z\}$ is a feedback vertex set of the graph $G$ because the difference between the graph $G\setminus F$ and the graph $G^{\prime}\setminus F^{\prime}$ is only an isolated vertex. Thus, $\textrm{fvs}(G)\leq\textrm{fvs}(G^{\prime})+2\leq 2k$.
@@
If there is a vertex $u$ of degree two in the graph $G$, which is not in a triangle, then let $G^{\prime}$ be the graph obtained by omitting the vertex $u$ from the graph $G$ and connecting the two neighbors of the vertex $u$ with an edge. This graph is trivially still a based planar graph. Any minimum feedback vertex set in this graph is also a minimum feedback vertex set in the original graph. If a minimum feedback vertex set in $G$ does not contain $u$, then it is a minimum feedback vertex set in $G^{\prime}$, too. If a minimum feedback vertex set in $G$ contains $u$, then let us delete $u$ from it and add a neighbor of $u$. So, we get a minimum feedback vertex set with the same size in $G^{\prime}$. Furthermore, similar is true for cycle packings: they are essentially identical in the two graphs. So $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, thus, by induction, we are done. (The statement is true for the graph $G^{\prime}$, so it is also true for the graph $G$.)
@@
Now, we may assume that the degree of each vertex in the graph is at least three. Then, according to Lemma 2.3, there is a triangle xyzxyz, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face. Let $G^{\prime}$ be the graph obtained by deleting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})<\textrm{cp}(G)$, since the vertex $x$ cannot be included in any cycle in the graph $G^{\prime}$, so we can add the triangle xyzxyz to an arbitrary cycle packing of the graph $G^{\prime}$ and thus get a bigger cycle packing. Then let us take a minimum feedback vertex set of the graph $G^{\prime}$, adding the vertices $y$ and $z$ to this, we get a feedback vertex set of the graph $G$, whose size is at most $2\textrm{cp}(G^{\prime})+2\leq 2\cdot\textrm{cp}(G)$. ∎
The proof works in the same way, if instead of cycle packing, we considered face packing, so the following stronger statement also holds.
Remark 2.5.
@@ 3 Concluding remarks
Conjecture 3.1.
Let $G$ be a planar graph in which there is a cycle which has at least one edge in common with every face. Prove $\textrm{fvs}(G)\leq 2\cdot\textrm{cp}(G)$.
@@
Let $G$ be a hamiltonian planar graph. Prove $\textrm{fvs}(G)\leq 2\cdot\textrm{cp}(G)$.